OpenStudy (ammarah):

Divide using polynomial synthetic division. (6x^3-7x+2) / (x-2)

OpenStudy (anonymous):

list the coefficients, don't forget to keep the place for the missing \(x^2\) term

OpenStudy (anonymous):

the coefficients are 6 0 -7 2

OpenStudy (anonymous):

then put a 2 on the side 6 0 -7 2 2_________________

OpenStudy (ammarah):

yeah i got 6x^2+12x+17+36/2 but i didnt get a zero so im confused...

OpenStudy (anonymous):

bring down the 6 6 0 -7 2 2_________________ 6

OpenStudy (anonymous):

oh ok no reason to necessarily get no remainder usually when you divide, there is a remainder, just like with numbers

OpenStudy (ammarah):

so my answer is correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well almost the remainder is not \(\frac{36}{2}\) it is \(\frac{36}{x-2}\)

OpenStudy (ammarah):

oh yes

OpenStudy (ammarah):

could u please help me with this: Find the product: (2x+3y)^3

OpenStudy (anonymous):

you only get zero if the polynomial factors exactly you should not think you are wrong if you do not get a zero as a remainder

OpenStudy (anonymous):

\[ (2x+3y)^3=(2x+3y)(2x+3y)(2x+3y)\] you can either grind it till you find it, or use \[(a+b)^3=a^3+3a^2b+3ab^2+b^3\] with \[a=2x,b=3y\]

OpenStudy (ammarah):

wait so can i use foil?

OpenStudy (ammarah):